The Knox Valley investment club invested $8000 at 4.6% annual simple interest. How much additional money should they invest at an annual simple interest rate of 3.6%, so that the total annual interest earned from the two accounts is 4.0% of the total investment?

Answer by stanbon(75887)

The Knox Valley investment club invested $8000 at 4.6% annual simple interest. How much additional money should they invest at an annual simple interest rate of 3.6%, so that the total annual interest earned from the two accounts is 4.0% of the total investment?
--------------
Amt.invested at 4.6% = 8000; interest on that is 0.046(8000) = 368 dollars
Amt.invested at 3.6% = x ; interest on that is 0.036x dollars
Amt. invested at 4% = 8000+x; interest on that is 0.04(8000+x) = 320+0.04x dollars
-------------------
EQUATIOn:
interest + interest = interest
368 + 0.036x = 0.040(8000+x)
368 + 0.036x = 320+0.040x
0.004x = 48
x = 12000 dollars (amt. of addtional money invested at 3.6%)
==========
Cheers,
Stan H.
EQUATION: